The polarisability of a given material is 1.0×10^-40 C^2 mN^−1 . Determine its dielectric constant if the number of molecules in 1 m^3 volume of the material is 5.0×10^25.
1
Expert's answer
2019-08-28T09:10:37-0400

According to the Clausius-Mosotti equation


"\frac{\epsilon-1}{\epsilon+2} = 1.33\pi N\alpha = 1.33\cdot 3.14\cdot 5\cdot10^{25}10^{-37}"


then "\epsilon" "=1 + 3\cdot 2.110^{-11} = 1+ 6.3 \cdot 10^{-11}"